# MRAG: A Modular Retrieval Framework for Time-Sensitive Question Answering

### v1

@@ Abstract
Understanding temporal relations and answering time-sensitive questions is crucial yet a challenging task for question-answering systems powered by large language models (LLMs). Existing approaches either update the parametric knowledge of LLMs with new facts, which is resource-intensive and often impractical, or integrate LLMs with external knowledge retrieval (i.e., retrieval-augmented generation). However, off-the-shelf retrievers often struggle to identify relevant documents that require intensive temporal reasoning. To systematically study time-sensitive question answering, we introduce the TempRAGEval benchmark, which repurposes existing datasets by incorporating temporal perturbations and gold evidence labels. As anticipated, all existing retrieval methods struggle with these temporal reasoning-intensive questions. We further propose Modular Retrieval (MRAG), a trainless framework that includes three modules: (1) Question Processing that decomposes question into a main content and a temporal constraint; (2) Retrieval and Summarization that retrieves evidence and uses LLMs to summarize according to the main content; (3) Semantic-Temporal Hybrid Ranking that scores each evidence summarization based on both semantic and temporal relevance. On TempRAGEval, MRAG significantly outperforms baseline retrievers in retrieval performance, leading to further improvements in final answer accuracy.
@@ 1 Introduction
Facts are constantly evolving in our ever-changing world. This dynamic nature highlights the need for natural language processing (NLP) systems capable of adapting to new information (Liška et al., 2022; Zhang et al., 2024; Kasai et al., 2024) and providing accurate responses to time-sensitive questions (Chen et al., 2021; Chu et al., 2024). For instance, a common query like “Who is the UK Prime Minister?” sees the answer transition from “Rishi Sunak” to “Keir Starmer” in 2024 (Figure 1).
@@
With developments of large language models (LLMs), existing approaches rely on the parametric knowledge of LLMs to answer time-sensitive questions directly, and constantly update the parametric knowledge on new facts (Rozner et al., 2024; Wu et al., 2024b; Wang et al., 2024). However, updating LLM parameters are often resource-intensive. Moreover, recent studies reveal that even after updates, LLMs often struggle to accurately reflect the latest facts and their associated facts (Cohen et al., 2024). An alternative line of research explores Retrieval-Augmented Generation (RAG), which integrates LLMs with external knowledge (e.g., Wikipedia) through information retrieval (Izacard et al., 2020). While RAG allows for the incorporation of new facts with minimal effort, its performance heavily relies on off-the-shelf retrieval systems, which are often limited to keywords or semantic matching. Time-sensitive questions, however, often require intensive temporal reasoning to identify relevant documents, i.e., reasoning-intensive retrieval (Su et al., 2024a). For example, in Figure 1, retrievers should infer the date “24 July 2019” as relevant to the constraint “as of 6 May 2021”, rather than only match with date like “May 2021”. Despite widespread recognition of the temporal reasoning-intensive retrieval challenge, there remains a lack of measurements and approaches. We fill that lacuna.
@@
We begin by conducting a diagnostic evaluation of existing retrieval approaches for temporal reasoning-intensive retrieval. Following the idea of systematic evaluation with contrast set (Gardner et al., 2020), we repurpose two existing datasets, TimeQA (Chen et al., 2021) and SituatedQA (Zhang and Choi, 2021), to introduce the Temporal QA for RAG Evaluation benchmark (TempRAGEval). We manually augment the test questions with temporal perturbations (e.g., modifying the time period). In addition, we annotate gold evidence on Wikipedia for more accurate retrieval evaluation. As expected, current retrieval methods exhibit limited temporal reasoning capabilities, especially on temporal perturbed questions.
@@
To address time-sensitive questions, we propose a training-free Modular Retrieval framework (MRAG) to enhance temporal reasoning-intensive retrieval. MRAG contains three key modules: (1) Question Processing, which decomposes each question into a main content and a temporal constraint; (2) Retrieval and Summarization, which adopts off-the-shelf retrievers to find evidence passages based on the main content, then instructs LLM to condense or split relevant passages into query-related sentences; (3) Semantic-Temporal Hybrid Ranking, which ranks each evidence sentence using a combination of a semantic score measuring semantic similarity, and a temporal score, a novel symbolic component that assesses temporal relevance to the query’s temporal constraint.
Compared to the classic retrieve-then-rerank systems, MRAG offers several advancements: (1) It disentangles relevance-based retrieval from temporal reasoning through question preprocessing, avoiding distractions of off-the-shelf retrievers; (2) It introduces a temporal scoring mechanism that evaluates whether the temporal information in the evidence (e.g.,2019 – 2022”) satisfies the query’s temporal relations (e.g.,before 2023”); (3) The modular framework enhances transparency, enabling users to easily identify which component leads to an incorrect answer.
@@
  • We introduce TempRAGEval, a time-sensitive QA benchmark dataset to diagnostically evaluate each component of existing retrieval-augmented generation systems.
  • We propose MRAG, a new modular framework to identify evidence through separating relevance-based retrieval and temporal reasoning.
  • On TempRAGEval, MRAG significantly outperforms all baseline retrieval systems, and the improvements further lead to better answer predictions.
@@ 2 Background
2.1 Time-sensitive Question Answering
Our task is to predict an answer $y$, to a time-sensitive question $q$, in which the input question $q$ also includes temporal relations $t_{1},...,t_{n}$. [^1] To simplify our setting, we make sure each question has a valid answer.
With advancements in LLMs, a straightforward approach is to directly ask these questions to an off-the-shelf LLM (e.g., LLama 3.1), hoping the relevant facts are captured in its parametric knowledge. However, this is often not the case for time-sensitive questions. A significant portion of such questions are recent events (e.g., “Who is the UK Prime Minister in November 2024?”), which typically fall beyond the knowledge cutoff for training LLMs (e.g., December 2023 for LLama 3.1).
One potential solution is to constantly update the parameters of LLMs with new facts, i.e., knowledge editing (Wu et al., 2024b; Wang et al., 2024). However, this approach is highly resource-intensive and often impractical for very large LLMs (e.g., GPT-4o). Moreover, knowledge editing is not always effective. Recent studies reveal that even after parameter updates, LLMs often fail to accurately answer questions about new facts (Liška et al., 2022; Rozner et al., 2024; Wang et al., 2024), let alone related or associated facts (Zhong et al., 2023; Wu et al., 2024b, c).
Therefore, our study adopts an alternative approach that leverages external text collections as a knowledge source, integrating relevant information into existing LLMs to answer questions (detailed in §2.2). While we use Wikipedia222we use December 2021 Wikipedia dump, with 33.1M text chunks as the text collection $D$, our approach is broadly applicable to any corpus, e.g., ClueWeb (Overwijk et al., 2022).
@@ 2.2 Retrieval-Augmented Generation
The goal of RAG is to address the limitations in the parametric knowledge of LLMs (Izacard et al., 2020). RAG involves a retrieval system to find relevant evidence from a large corpus and a LLM to generate a response based on the evidence (Wu et al., 2024a).
@@
Retrieval methods are typically categorized into sparse retrieval and dense retrieval. Sparse retrieval methods, like TF-IDF (Das and Chakraborty, 2018) and BM25 (MacAvaney et al., 2020), rely on lexical matching. In contrast, dense retrieval models (Karpukhin et al., 2020; Zhao et al., 2021; Thakur et al., 2021; Izacard et al., 2021) use LLMs (e.g., BERT) encode the question $q$ and the passage $p$ separately using two independent encoders (i.e., bi-encoders). These models learn a scoring function (e.g., dot product) between question and passage vectors to capture semantic similarities:
@@
Dense retrieval models are highly scalable, since passages can be encoded offline, and are efficiently retrieved over maximum inner product search (MIPS) with the question (Shrivastava and Li, 2014).
However, bi-encoder models may lack the ability to capture fine-grained interactions between the query and passage. A common optional [^3] Note that reranking is not always adopted, as it adds additional computational cost. approach is to have another cross-encoder model to rerank top retrieved passages. Cross-encoder models (Khattab and Zaharia, 2020; Wang et al., 2020; Gemma et al., 2024) jointly encode the query $q$ and the passage $p$ together by concatenating them as input into a single model as:
@@
The answer generation component takes the question and support passages as input to generate the answer. A widely used approach is Fusion-in-Decoder (Izacard et al., 2020), which fine-tunes an encoder-decoder LLM (e.g., T5 (Raffel et al., 2020)). FiD independently encodes each passage and question, then concatenates their representations into the decoder, as formulated below:
$y=\text{Dec}([\text{Enc}([q;p_{1}]);\dots;\text{Enc}([q;p_{k}])]),p_{k}\in\mathcal{D}.$ (3)
With decoder-only LLMs (OpenAI et al., 2024; Dubey et al., 2024), the support passages and the question are concatenated into a single sequence.
To unlock the reasoning capabilities of LLMs, Chain-of-Thought (CoT) prompting (Wei et al., 2022) introduces intermediate reasoning steps within prompt examples, significantly improving performance on various reasoning tasks. Self-RAG (Asai et al., 2024) predicts answers for each passage independently, and then adopts self-reflect mechanisms (Yao et al., 2022) into prompts to select the best answer. Self-RAG achieve SOTA results on multiple knowledge-intensive QA datasets.
@@ 3 TempRAGEval Benchmark
3.1 Existing Time-Sensitive QA Datasets
Existing time-sensitive QA datasets often feature questions where timestamps are closely tied to their answers, as illustrated in Figure 1. Consequently, retrievers relying on surface-form date matching tend to perform well on these test sets, limiting the evaluation of true temporal reasoning capabilities. To systematically evaluate time-sensitive question answering, we establish the following criteria: (1) questions with varying difficulty levels, including both common and rare entities; (2) manual evidence annotation to enable retrieval evaluation; and (3) temporal perturbations as contrast sets (Gardner et al., 2020). Since no existing datasets provide both temporal perturbations and evidence annotations, we primarily adapt from the following two datasets:
  • SituatedQA (Zhang and Choi, 2021) is a open-domain QA dataset where the correct answer to a question is based on a temporal context, specifically of the type “as of”. SituatedQA sources questions from exiting QA datasets such as NQ-Open (Kwiatkowski et al., 2019) and WebQuestions (Berant et al., 2013), which primarily focus on popular entities.
  • TimeQA (Chen et al., 2021) is a time-sensitive QA dataset in which question-answer pairs are synthetically generated from time-evolving WikiData facts using templates. The dataset includes a wide range of reasoning types and difficulty levels, with a notable emphasis on rare entities.
Other time-sensitive QA datasets, StreamingQA Liška et al. (2022), TempLAMA (Dhingra et al., 2022), and concurrent dataset ComplexTQA (Gruber et al., 2024), can also serve as alternatives. Temporal reasoning benchmarks such as Wei et al. (2023); Chu et al. (2024); Wang and Zhao (2024); Tan et al. (2023) fall outside the scope of our study.
3.2 TempRAGEval Construction
Dataset # Eval. Evid. Natu. Pert.
ComplexTQA 10M
@@ 3.2 TempRAGEval Construction
TimeQA 3K
MenatQA 2K
@@
Table 1: Comparison of temporal QA datasets. TempRAGEval is featured by manual evidence annotations, human-written question (i.e., Naturalness), and perturbations.
@@
We source questions from TimeQA (Chen et al., 2021) and SituatedQA (Zhang and Choi, 2021) datasets. For each question, we introduce perturbations to the original temporal constraints, e.g.,as of 2019”, with complex variations such as e.g.,between 2020 and 2021”, “the last before 2020”, or “as of 6 May 2021”. We guarantee the answer (e.g.,Boris Johnson”) remains consistent across the perturbed questions.
@@
To better evaluate the performance of retrieval systems, we supplement question-answer pairs with up to two annotated gold evidence passages. Specifically, for each question, we manually review top-100 passages retrieved by the leading retriever, Contriever (Izacard et al., 2021), and the best reranker, Gemma (Li et al., 2023). If these retrieval methods fail to identify relevant passages within top-100 passages, we manually search Wikipedia pages related to the query entities to locate the gold evidence (around 12.7% of original questions). In total, we have annotated 1,000 test examples with both original and perturbed questions, along with human-annotated evidence. TempRAGEval-TimeQA (500 examples) focus on popular entities, and TempRAGEval-TimeQA (500 examples) on rare entities.
@@ 3.3 Preliminary Evaluation on TempRAGEval
Figure 2: The retrieval performance degradation of the Gemma baseline on TempRAGEval-SituatedQA, comparing original and perturbed questions (see TempRAGEval-TimeQA in Appendix C).
@@ 4 MRAG: Modular Retrieval
We apply off-the-shelf retrievers (e.g., Contriever) to find relevant passages to MC in Wikipedia. Then we adopt a cross-encoder model to rerank the passages with high semantic similarity to MC.
It is common for a passage to contain multiple pieces of temporal information, most of which are unrelated to the question and can distract the temporal scoring component introduced next. For example, the relevant passage in Figure 3 includes the sentence, “it was filmed in 2017”, which satisfies the TC but is irrelevant. Therefore, we initially considered splitting passages into individual sentences to eliminate temporal distractors. However, as shown in Figure 3, critical information from the most relevant passages—such as “America’s Next Top Model”, “The winner of the competition”, and ”January 9, 2018”—is often scattered across different sentences. Relying solely on sentence splitting would miss key details. To overcome this challenge, we instead adopt LLMs to summarize the top-k passages into a single sentence to condense relevant phrases and temporal information. [^4] While using LLMs to summarize passages can introduce hallucinations, we mitigate this by summarizing only the top-k passages and simply splitting the remaining passages into sentences. An analysis of this approach is provided in §6.1.
@@
We rerank each sentence (summarized from a LLM or segmented from the original passage) with two distinct scores: a semantic score and a temporal score. The semantic score is calculated from the similarity between the evidence sentence and MC. For temporal score, we first extract temporal information in the evidence sentence, then we apply symbolic scoring functions [^5] We include details of pre-defined temporal scoring functions in Appendix D. to verify whether the temporal information in the evidence satisfies TC. For example, the evidence summary timestamp “2018” satisfies TC “as of 2021” in Figure 3. The final score for each sentence is the product of the semantic score and the temporal score. We select the passages that contains the highest-scoring sentences.

### v2

@@ Abstract
Understanding temporal concepts and answering time-sensitive questions is crucial yet a challenging task for question-answering systems powered by large language models (LLMs). Existing approaches either update the parametric knowledge of LLMs with new facts, which is resource-intensive and often impractical, or integrate LLMs with external knowledge retrieval (i.e., retrieval-augmented generation). However, off-the-shelf retrievers often struggle to identify relevant documents that require intensive temporal reasoning. To systematically study time-sensitive question answering, we introduce the TempRAGEval benchmark, which repurposes existing datasets by incorporating complex temporal perturbations and gold evidence labels. As anticipated, all existing retrieval methods struggle with these temporal reasoning-intensive questions. We further propose Modular Retrieval (MRAG), a trainless framework that includes three modules: (1) Question Processing that decomposes question into a main content and a temporal constraint; (2) Retrieval and Summarization that retrieves, splits, and summarize evidence passages based on the main content; (3) Semantic-Temporal Hybrid Ranking that scores semantic and temporal relevance separately for each fine-grained evidence. On TempRAGEval, MRAG significantly outperforms baseline retrievers in retrieval performance, leading to further improvements in final answer accuracy. [^1] Our code and data are available at https://github/anonymous.
@@ 1 Introduction
Facts are constantly evolving in our ever-changing world. This dynamic nature highlights the need for natural language processing (NLP) systems capable of updating information (Liška et al., 2022; Zhang et al., 2024; Kasai et al., 2024) and providing accurate responses to time-sensitive questions (Chen et al., 2021; Chu et al., 2024). For instance, a common query like “Who is the UK Prime Minister?” sees the answer transition from “Boris Johnson” to “Rishi Sunak” in 2022 (Figure 1).
@@
With developments of large language models (LLMs), existing approaches rely on the parametric knowledge of LLMs to answer time-sensitive questions directly, and constantly update the parametric knowledge on new facts (Rozner et al., 2024; Wu et al., 2024a; Wang et al., 2024). However, updating LLM parameters are often resource-intensive. An alternative line of research explores Retrieval-Augmented Generation (RAG), which integrates LLMs with external knowledge (e.g., Wikipedia) through information retrieval (Izacard et al., 2020). While RAG allows for the incorporation of new facts with minimal effort, its performance heavily relies on off-the-shelf retrieval systems, which are often limited to keywords or semantic matching. Time-sensitive questions, however, often require intensive temporal reasoning to identify relevant documents, i.e., reasoning-intensive retrieval (Su et al., 2024a). For example, in Figure 1, retrievers should infer the date “24 July 2019” as relevant to the constraint “as of 6 May 2021”, rather than only match with date like “May 2021”. Although temporal QA is a well-established domain, there remains a lack of research for temporal reasoning-intensive retrieval systems. We fill that lacuna.
@@
We begin by conducting a diagnostic evaluation of existing retrieval approaches for temporal reasoning-intensive retrieval. Following the idea of systematic evaluation with contrast set (Gardner et al., 2020), we repurpose two existing datasets, TimeQA (Chen et al., 2021) and SituatedQA (Zhang and Choi, 2021), to introduce the Temporal QA for RAG Evaluation benchmark (TempRAGEval). We manually augment the test questions with complex temporal perturbations (e.g., modifying the time period to avoid textual overlap). In addition, we annotate gold evidence on Wikipedia for more accurate retrieval evaluation. We observe remarkable degradation of retrieval performance, which points out a serious robustness issue for existing retrievers.
@@
To address this issue, we propose a training-free Modular Retrieval framework (MRAG) to enhance temporal reasoning-intensive retrieval. Complementary to recent agentic RAG methods such as R1-Searcher and Search-o1 (Song et al., 2025; Li et al., 2025), MRAG provides an improved retrieval component and reduces the number of invocations for retrieval when handling complex temporal questions. Specifically, MRAG contains three key modules: (1) Question Processing, which decomposes each question into a main content and a temporal constraint; (2) Retrieval and Summarization, which utilizes off-the-shelf retrievers to find evidence passages based on the main content, segments them into independent sentences, and guides LLMs to condense the most relevant passages into query-specific sentences.(3) Semantic-Temporal Hybrid Ranking, which ranks each evidence sentence using a combination of a semantic score measuring semantic similarity, and a temporal score, a novel symbolic component that assesses temporal relevance to the query’s temporal constraint.
@@
  • We introduce TempRAGEval, a time-sensitive RAG benchmark to diagnostically evaluate each component of existing retrieval-augmented generation systems.
  • We propose MRAG, a modular retrieval framework to separately determine semantic and temporal relevance.
  • On TempRAGEval, MRAG significantly outperforms all baseline retrieval systems, and the improvements lead to better answer generation.
@@ 2 Background
2.1 Temporal Question Answering
There has been extensive research on Temporal Question Answering over Knowledge Graphs (TKGQA). Prior works have developed knowledge graph-based benchmarks (e.g., CronQuestions (Saxena et al., 2021) and TimeQuestions (Jia et al., 2021)), graph neural network models (Chen et al., 2022; Liu et al., 2023; Sharma et al., 2023), and TKG-augmented LLMs (Qian et al., 2024). However, these methods heavily rely on knowledge graphs, which are costly to build, domain-specific, and prone to becoming outdated (Kau et al., 2024). Therefore, we focus on the free-text setting, which is more universal and up-to-date. Instead of building temporal-specific graph models, we improve general-purpose text retrievers for temporal queries. We use Wikipedia as the main knowledge corpus $D$ [^2] December 2021 Wikipedia dump, comprising 33.1 million text chunks. Our approach is broadly applicable to other collections, such as the New York Times Annotated Archive (Sandhaus, 2008) and ClueWeb (Overwijk et al., 2022).
@@ 2.2 Retrieval-Augmented Generation
The goal of RAG is to address the limitations in the parametric knowledge of LLMs by incorporating external knowledge.
@@
Retrieval methods are typically categorized into sparse retrieval and dense retrieval. Sparse retrieval methods like BM25 (MacAvaney et al., 2020) rely on lexical matching. In contrast, dense retrieval models (Karpukhin et al., 2020; Zhao et al., 2021; Thakur et al., 2021; Izacard et al., 2021) encode the question $q$ and passage $p$ into low-dimension vectors separately. The semantic similarity is computed using a scoring function (e.g., dot product) as:
@@
However, these bi-encoder models lack the ability to capture fine-grained interactions between the query and passage. A common optional [^3] Note that reranking is not always adopted, as it adds additional computational cost. approach is to have another cross-encoder model to rerank top passages. Cross-encoder models (Khattab and Zaharia, 2020; Wang et al., 2020; Gemma et al., 2024) jointly encode the query $q$ and the passage $p$ together by concatenating them as input into a single model as:
@@
Recent reader systems are mainly powered by LLMs with strong reasoning capabilities. With recent advancements in long-context LLMs (Dubey et al., 2024; Lee et al., 2024b), top documents are typically concatenated with the query as reader input:
$y=\text{Dec}(\text{Enc}([p_{1};\dots;p_{k};q])).$ (3)
To unlock the reasoning capabilities of LLMs, Chain-of-Thought (CoT) prompting (Wei et al., 2022) introduces intermediate reasoning steps for improved performance. Self-RAG (Asai et al., 2024) critiques the retrieved passages and its own generations. Recent agentic RAG systems dynamically acquire and integrate external knowledge (using either retriever models or search APIs) during the reasoning process (Li et al., 2025; Song et al., 2025; Jin et al., 2025).
@@ 3 TempRAGEval Benchmark
Dataset # Eval. Evid. Natu. Comp.
ComplexTQA 10M $\checkmark$
@@
TimeQA 3K $\checkmark$
MenatQA 2K $\checkmark$
@@
Table 1: Comparison of temporal QA datasets. TempRAGEval is featured by manual evidence annotations, human-written question (i.e., Naturalness), and higher complexity in temporal reasoning.
3.1 Existing Time-Sensitive QA Datasets
There are several existing QA datasets that focus on temporal reasoning. The most representative ones are the following: [^4] We mainly focus on these two datasets, while others, such as Liška et al. (2022); Dhingra et al. (2022); Gruber et al. (2024), serve as alternatives. Datasets such as Tan et al. (2023); Virgo et al. (2022) that are not knowledge-intensive, are excluded from this work (Appendix B).
  • SituatedQA (Zhang and Choi, 2021) is a time-sensitive QA dataset where the answer to an information-seeking question varies based on temporal context. These questions contain a single type of temporal constraint (e.g.,as of”) that directly align with the answers. Retrievers with surface-form date matching often exploit these shortcuts to bypass the need for temporal reasoning.
  • TimeQA (Chen et al., 2021) is another time-sensitive QA dataset. Unlike SituatedQA, the questions in hard split include complex temporal constraints (e.g.,between 2012 to 2018”). However, question-answer pairs are synthetically generated from time-evolving WikiData facts using templates. In addition, TimeQA does not include evidence annotations, making it imprecise to evaluate retrieval results.
According to Table 1, we observe that none of existing datasets include these key factors for systematically evaluating current retrieval (and answer generation) systems: (1) Evidence annotation; (2) Natural questions from users; (3) Complex temporal constraints. Therefore, as we will show in the following section, we aim to address this limitation by creating TempRAGEval.
3.2 TempRAGEval Construction
@@
Annotators first select question-answer pairs from the SituatedQA and TimeQA datasets [^5] Since both datasets lack questions about knowledge beyond the cutoff date of existing LLMs, we primarily focus on historical knowledge and discuss potential future directions on recent knowledge in the Limitations section. that can be grounded in Wikipedia facts with key timestamps or durations. They then revise temporal perturbations by selecting implicit conditions, temporal relations, and alternative dates to include complex temporal reasoning, without changing the final answer [^6] Questions with the same content but different temporal constraints and answers are considered different samples. Perturbations are introduced for each sample.. Annotators are also required to edit the question text to improve naturalness. We include detailed guidelines in Appendix C.
@@
To better evaluate the performance of retrieval systems, we supplement question-answer pairs with up to two annotated gold evidence passages. A passage is relevant to the question if annotators can obtain the correct answer based on the passage. Specifically, for each question, annotators are asked to manually review top-20 passages retrieved by Contriever (Izacard et al., 2021) and reranked by the best Gemma (Li et al., 2023) reranker. If there is no relevant passage, annotators are required to search Wikipedia pages related to the query entities to locate the gold evidence (around 12.7% of questions). We create 1,000 test examples with human-annotated evidence. Appendix D presents sample statistics, revealing that SituatedQA questions include popular entities while the entities for TimeQA questions are long-tailed.
@@ 3.3 Preliminary Evaluation on TempRAGEval
Figure 2: The retrieval performance degradation of the Gemma baseline on TempRAGEval-SituatedQA, comparing original and perturbed questions (see TempRAGEval-TimeQA in Appendix E).
@@ 4 MRAG: Modular Retrieval
We apply off-the-shelf retrievers (e.g., Contriever) to find relevant passages to MC in Wikipedia. Then we employ reranker models to reorder these passages by semantic similarity to MC.
It is common for a passage to contain multiple pieces of temporal information, most of which are unrelated to the question and can distract the temporal scoring component introduced next. For example, the relevant passage in Figure 3 includes the sentence, “it was filmed in 2017”, which satisfies the TC but is irrelevant. Therefore, we split passages into individual sentences to eliminate temporal distractors. However, as shown in Figure 3, critical information from the most relevant passages—such as “America’s Next Top Model”, “The winner of the competition”, and ”January 9, 2018”—can be scattered across different sentences. Relying solely on sentence splitting would miss key details. To overcome this challenge, we additionally employ LLMs to summarize each of the top-k passages into a single sentence condensing relevant phrases and temporal information, as analyzed in §6.1.777While using LLMs to summarize passages can introduce hallucinations, we mitigate this by summarizing only the top-k passages.
@@
We rerank each sentence (summarized from a LLM or segmented from the original passage) with two distinct scores: a semantic score and a temporal score. The semantic score is calculated from the similarity between the evidence sentence and MC. For temporal score, we first extract the timestamp from each sentence (e.g.,2018”). Based on the timestamp and TC (e.g.,as of 2021”), we compute a temporal score using symbolic functions similar to temporal activation functions in Chen et al. (2022). The final score for each sentence is obtained by multiplying the semantic score and the temporal score. Finally, we select the passages that contains the highest-scoring sentences. We include the details of symbolic functions in Appendix G.